OpenStudy (anonymous):

HELP! How to find the inverse of f(x) = x - 2

OpenStudy (anonymous):

@satellite73 @jim_thompson5910 @texaschic101 @myininaya @zepdrix

OpenStudy (texaschic101):

f(x) = x - 2......same as y = x - 2 y = x - 2.....switch the x and y x = y - 2 -- now solve for y x + 2 = y or y = x + 2 so inverse would be : f ^-1(x) = x + 2 http://coolmath.com/algebra/16-inverse-functions/05-how-to-find-the-inverse-of-a-function-01.htm